is (8,-3) a solution of 4x+5y=-5 and what are the steps to do it?
Answer by checkley77(12844)

(8,-3)
4x+5y=-5
4*8+5*-3=-5
32-15=-5
17 DOES NOT = -5 .
THEREFORE: [X=8 & Y=-3 IS NOT A SOLUTION.]